OpenStudy (anonymous):

Simplify (-2)^2 + |-15| ÷ 3 11/3 19/3 9

OpenStudy (anonymous):

what is the value of \[\left| -15 \right|\]?

OpenStudy (anonymous):

@magan are you there?

OpenStudy (anonymous):

|-15| =+15 (-2)^2=4 We have : 4+15/3 First do 15/3 = 5 then have : 4+5=9 Got it ?:)

OpenStudy (anonymous):

only 15 we remove minus

OpenStudy (anonymous):

The purpose of this website is not to copy-paste answers from people that know the answer, but to help you solve the answers yourself. If you keep taking the easy way out, you will never truly understand. The best way of learning is by practicing.

OpenStudy (anonymous):

\( \large\bf You~can~read~the~\underline{\href{ /code-of-conduct }{Code~of~Conduct}}~here. \)